OpenStudy (anonymous):

In △PQR, find the measure of ∡P. 30.4° 35.9° 59.6° 54.1°

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Let us see where have you reached.. Just use again that sin formula, here you are given with perpendicular and hypotenuse, just find the angle here.. :)

OpenStudy (anonymous):

I think its 30.4?

OpenStudy (anonymous):

Not exactly 30 ??

OpenStudy (anonymous):

uhm...

OpenStudy (anonymous):

now i got 54.1

OpenStudy (anonymous):

how would you set it up

OpenStudy (anonymous):

cos(x) = 33.8/57.6

OpenStudy (anonymous):

Why cos(x) ??

OpenStudy (anonymous):

Perpendicular/hypotenuse is sin(x), I have earlier told you, remember it..

OpenStudy (anonymous):

;/

OpenStudy (anonymous):

Now find it..

OpenStudy (anonymous):

so then how would it be set up:( I'm confused

OpenStudy (anonymous):

i got .5868

OpenStudy (anonymous):

You have set it up correctly, but in place of cos(x), use sin(x) and find x then..

OpenStudy (anonymous):

How you are getting that??

OpenStudy (anonymous):

33.8/57.6?

OpenStudy (anonymous):

Yes, this is right.. But on Left hand side why are you using cos(x)??

OpenStudy (anonymous):

i give up, i rlly do not understand anything!

OpenStudy (anonymous):

I am talking about: \[\color{red}{\cos(x)} = \frac{33.8}{57.6}\] Are you seeing what I have written in red above?? There, write sin(x)

OpenStudy (anonymous):

sin^-1 (33.8/57.6) and i got 35.9?

OpenStudy (anonymous):

Yes, this is right but you are not getting what you are doing, you are just doing them not understanding them.. :)

OpenStudy (anonymous):

No matter how many of them you do, you are getting nothing.. :P

OpenStudy (anonymous):

? would that be the answer...

OpenStudy (anonymous):

35.9..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

What now??

OpenStudy (anonymous):

ok, theres no reason to be rude. Im asking for your help....

OpenStudy (anonymous):

I am not getting that what you have said.. I have written your answer above and you are still asking it again.. :P

OpenStudy (anonymous):

Okay, I am sorry if I sounded like that.. :(

OpenStudy (anonymous):

ok. i did not know that was the correct answer because you wrote (...) so just say yes or no.

OpenStudy (anonymous):

its ok